A cube has an edge of 2x. A larger cube has an edge of 4x. What is the ratio of the volume of the larger cube to the volume of the smaller cube?
A) 1:2 B) 1:8 C) 2:1 D) 4:1 E) 8:1

Answer by richwmiller(17219)

(2*2x)^3/(2x)^3=8/1=8:1